OpenStudy (anonymous):

For questions 5 and 6, recall that, when interest is compounded continuously, the balance in an account after t years is given by A = Pe^rt, where P is the initial investment and r is the interest rate. 5. How long will it take for $2000 to double if it is invested at 6.25% interest compounded continuously? 6. What rate of interest compounded continuously is needed for an investment of $500 to grow to $900 in 10 years?

OpenStudy (anonymous):

problem 5, solve \(e^{.065t}=2\) for t in two steps a) \(.0625t=\ln(2)\) b) \( t=\frac{\ln(2)}{.0625}\)

OpenStudy (anonymous):

second one solve \(900=500e^{10r}\) for r , steps are very similar

OpenStudy (anonymous):

divide by 500 get \[1.8=e^{10r}\] \[\ln(1.8)=10r\] \[r=\frac{\ln(1.8)}{10}\]

OpenStudy (anonymous):

That all looks great but I'm so confused.. I need to find a resource to help explain it to me..Online classes are so tough!

OpenStudy (anonymous):

So would the answer be 0.05877?

OpenStudy (anonymous):

I got the first one figured out. :) Just double checking the second one